OpenStudy (anonymous):

Implicit differentiation: 3(x^2+y^2)^2=100xy

OpenStudy (anonymous):

\[3(x ^{2}+y ^{2})^{2}=100xy\] How should I go about differentiating this? Multiply everything out, or just dive straight in and use the chain rule?

OpenStudy (anonymous):

\[6(x^2+y^2)(2x+2y \frac{dy}{dx})=100(y+x \frac{dy}{dx})\]

OpenStudy (anonymous):

So you're using the chain rule, then just expand it out, then solve? Also, why is "(x)(dy/dx)" on the other side of your "="